OpenStudy (anonymous):

rationalizing denominators, the cube root of 2y squared divided by the cube root of 9x squared

OpenStudy (anonymous):

\[\frac{\sqrt[3]{2y}}{\sqrt[3]{9x^2}} \times \frac{\sqrt[3]{9x^2}}{\sqrt[3]{9x^2}}\]\[\implies\frac{\sqrt[3]{2y * 9x^2}}{9x^2}\] \[\implies\frac{\sqrt[3]{18yx^2}}{9x^2}\]

OpenStudy (amistre64):

cbrt(9x^2)^2 does not equal 9x^2...

OpenStudy (amistre64):

cbrt((3^2 x^2) times (cbrt(3x)) = 3x

OpenStudy (anonymous):

Sorry, I was thinking of a regular square root, thanks for catching that amistre. :P

OpenStudy (amistre64):

no prob; happens to me all the time :)

OpenStudy (amistre64):

trig test in 20mins...yawn :)

OpenStudy (anonymous):

So the denominator should be \[\sqrt[3]{81x^4}\].

OpenStudy (amistre64):

denom will rationalize to 3x

OpenStudy (amistre64):

cbrt(3^2 x^2 3 x) = cbrt(3^3 x^3) = 3x

OpenStudy (anonymous):

\[thats what I got but the answer \in the back of the book says cube \root of6xy squred divided by 3x\]

OpenStudy (anonymous):

You're right. :P So the numerator should, in fact, be \[\sqrt[3]{6yx}\] and the denominator\[3x.\]

OpenStudy (amistre64):

top: cbrt(2y) cbrt(3x) = cbrt(6xy) yep; checks out on my "mr proffesor"

OpenStudy (radar):

Is it\[\sqrt[3]{(2y)2}\] or\[\sqrt[3]2y{2}\]

OpenStudy (anonymous):

how?

OpenStudy (amistre64):

Ciao :)

OpenStudy (radar):

The first expression is sort of confusing as it is written.

OpenStudy (anonymous):

@abdon: If you multiply the denominator by cbrt(3x) you get cbrt(27*x^3) which simplifies to 3x. Now, because you multiplied the denominator by that factor, you have to multiply the numerator by the same factor, and you get cbrt(6xy).

OpenStudy (radar):

My first post was just seeking clarification.

OpenStudy (anonymous):

Ohhhhhhh, sorry, it is y^2...but that stays constant throughout, didn't see that before.

OpenStudy (radar):

I am talking about the original problem posted by abdon

OpenStudy (radar):

What is squared the 2y or just the y

OpenStudy (anonymous):

I believe it was \[\sqrt[3]{2y^2}/\sqrt[3]{9x^2}.\]

OpenStudy (radar):

OK. Thanks, now I will see if I get the same as you.

OpenStudy (anonymous):

Sorry for having such broken up responses, here's the condensed form of what I got:\[\frac{\sqrt[3]{6xy^2}}{3x}.\]

OpenStudy (anonymous):

\[got class ,thanks still confusing\]